# Group-On: Boosting One-Shot Segmentation with Supportive Query

### v1

@@ Appendix A Analysis of Attention Correlations for Group-On Voting
First, the low-level feature maps $\textbf{{F}}_{\text{low}}^{\text{h}},\textbf{{F}}_{\text{low}}^{\text{s}}\in\mathbb{R}^{H_{1}\times W_{1}\times C_{1}}$ of the host query and supportive query are extracted from the 3th layer of the total features. Then we reshape them and calculate the Gram Matrices, following the previous work[26]. Please note that the relevant operations of these two queries are similar, and that of the host one can be summarized as:
$\textbf{{A}}_{\text{h}}=\mathcal{F}_{\text{reshape}}(\textbf{{F}}_{\text{low}}^{\text{h}})\in\mathbb{R}^{C_{1}\times N_{1}},$ (11)
$\textbf{{G}}^{\text{h}}=\textbf{{A}}_{\text{h}}\textbf{{A}}_{\text{h}}^{\mathsf{T}}\in\mathbb{R}^{C_{1}\times C_{1}},$ (12)
@@
where $N_{1}=H_{1}\times W_{1}$ and $\mathcal{F}_{\text{reshape}}$ reshapes the size of input tensor to $C_{1}\times N_{1}$.
Appendix B Comparison of Using Different Query Amounts for Testing
To further investigate the impact of varying query amounts during testing, we carry out supplementary experiments using larger numbers of query images. In Fig. 7, as the number of query images varies from 1 to 30, the inference performance climbs up and remains relatively stable without any noticeable downtrends. However, the inference time per query manifests an sustainable growth. Such trends indicate a marginal performance improvement at the cost of significant computational resources after reaching a point of stable fluctuation. In light of the inference efficiency, it is not a viable strategy to increase the query amount without restrictions. Instead, one should select the most favorable query amount that balances both efficiency and performance. In this paper, we set the query amount to 5 during testing.
Figure 7: Supplementary study on the impact of query amounts for testing performed on fold-0 of the PASCAL-5i [5] dataset.

### v2

@@ A-D Interpretability of Router Weights in MoME
First, the low-level feature maps $\textit{F}_{\text{low}}^{\text{host}},~{}\textit{F}_{\text{low}}^{\text{supp}}\in\mathbb{R}^{H_{1}\times W_{1}\times C_{1}}$ of the host query and supportive query are extracted from the 3th layer of the total features. Then we reshape them and calculate the Gram Matrices, following the previous work[21]. Please note that the relevant operations of these two queries are similar, and that of the host one can be summarized as:
$\textit{A}_{\text{host}}=\mathcal{F}_{\text{reshape}}(\textit{F}_{\text{low}}^{\text{host}})\in\mathbb{R}^{C_{1}\times N_{1}},$ (8)
$\textit{G}^{\text{host}}=\textit{A}_{\text{host}}\textit{A}_{\text{host}}^{\mathsf{T}}\in\mathbb{R}^{C_{1}\times C_{1}},$ (9)
@@
where $N_{1}=H_{1}\times W_{1}$ and $\mathcal{F}_{\text{reshape}}$ reshapes the size of input tensor to $C_{1}\times N_{1}$. With the calculated Gram Matrices, the Frobenius norm is evaluated on their difference to obtain the overall indicator $\psi$:
$\psi=\|\textit{G}^{\text{host}}-\textit{G}^{\text{supp}}\|_{F},$ (10)
where $\|\cdot\|_{F}$ denotes the Frobenius norm of the input matrix.
A-E Qualitative Results
To intuitively display our approach, we visualize some results containing challenging intra-class variation under the one-shot setting in Fig. 7. The first row is support images with their masks in blue, while the second row represents the groundtruth. And the following two rows display predictions of the baseline model HSNet [6] and our approach, respectively. Note that all the images in each column belong to the same category. One can see from our results (the 4th row) that if objects in the support image differ greatly from those in the query images, our Group-On approach captures the semantic essence of the objects and segments more accurately with the reciprocal knowledge help of the multiple query images.